Question Hep * A pharmaceutical company receives large shipments of aspirin tablets. The acceptance batch if there is only one or none that doesnt meet the required specifications If one shipment of 7000 aspirin tablets probability that this whole shipment will be sampling plan is to randomly select and test 46 tablets, then accept the whole C actually has a 6% rate of defects, what is the or will many be rejected? The probability that this whole shipment will be accepted is (Round to four decimal places as needed ) ter your answer in the answer box and then click Check Answer
media%2F7ad%2F7adf5c3e-9267-4b27-b8d3-3f
Answer #1

1)P(shipment will be accepted)=P(X<=1)=P(X=0)+P(X=1)=46C0(0.06)0(0.94)46+46C1(0.06)1(0.94)45

=0.2285

2)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 105
std deviation   =σ= 15.0000
probability = P(88<X<122) = P(-1.13<Z<1.13)= 0.8708-0.1292= 0.7416